# Chirality induced phase separation in active circle swimmers

### v1

@@ Abstract
Many microswimmers are inherently chiral, and this chirality can introduce fascinating behaviors in a collection of microswimmers. The dynamics become even more intriguing when two types of microswimmers with distinct chirality are mixed. Our study considers a mixture of self-propelled particles with opposite chirality, examining how the system’s characteristics evolve as the magnitude of chirality is tuned. In weakly chiral systems, the particles exhibit similar behavior, leading to a globally flocking phase where both types of particles are well-mixed. However, in an intermediate range of chirality, the particles demix and follow their trajectories, creating a competition between chirality and self-propulsion. This competition results in interesting phases within the system. We explore the characteristics of these different phases in detail, focusing on the roles of self-propulsion speed and chirality.
@@ I Introduction
Active matter is a fascinating area of study Sumpter (2010); Vicsek and Zafeiris (2012); Marchetti et al. (2013); Lavi et al. (2024); Shankar et al. (2022) that focuses on self-propelled entities such as active colloidsBechinger et al. (2016), insects swarm in large groups Cavagna et al. (2017), fish display schooling in the ocean Ward et al. (2008), and birds flock in the sky Ballerini et al. (2008), groups of animals at the macroscopic scale Cavagna and Giardina (2014), etc. These entities move due to internal mechanisms such as chemical reactions, cilia, flagella, and internal motors utilizing the input from the surrounding mediums. Understanding collective behavior in active matter is crucial for uncovering the interpretation of collective behavior. Flocking, which refers to the collective motion of organisms, is also observed in inanimate macroscopic systems, such as active granular rods Kumar et al. (2014). This widespread phenomenon often involves transitions from disordered to long-range ordered states, influenced by factors like density and noise Vicsek et al. (1995).
@@
The dynamics of active agents can be categorized into two main types: one involves nearly straight movement with random directional changes, creating a sense of self-propulsion without any perception of rotation, as seen in active colloids like Janus particles, molecular motors Jülicher et al. (1997), natural low-Reynolds-number swimmers Najafi and Golestanian (2004); Ebrahimian et al. (2015); Bennett and Golestanian (2013), active particles on periodic lattice Chamolly et al. (2017) and polar active granular particles Zhang et al. (2017). The second type is characterized by particles, which have a sense of rotation or chiral particles. Systematically, particles can tend to rotate clockwise or anticlockwise. The motion is circular in two dimensions and helical in three dimensions Löwen (2016); Friedrich (2016). These chiral particles possess a characteristic self-propulsion velocity and a chirality that determines the rate of directional change in the absence of fluctuations. The radius of their trajectories is determined by the ratio of the self-propulsion velocity to the chirality. This type of chiral motion is observed in various systems, including biological circle swimmers like bacteria DiLuzio et al. (2005), E.coli Berg and Turner (1990); Lauga et al. (2006), sperm cells Riedel et al. (2005); Elgeti et al. (2015), L-shape artificial microswimmers Kümmel et al. (2013) and magnetotactic bacteria in rotating external fields Ērglis et al. (2007); Cēbers (2011).
@@
While most research on bacterial swarming and collective swimming has concentrated on homogeneous systems Harshey (2003); Sokolov et al. (2007); Copeland and Weibel (2009); Chatterjee et al. (2023), which involve agents having identical features. The natural world notably encounters mixtures of particles of different chiralities Schumacher et al. (2017); Jolles et al. (2020). These differences arise from variations in their size Peled et al. (2021), diverse orientations, and mixed-species behavior Benisty et al. (2015); Zuo and Wu (2020). Complex systems are inherently heterogeneous, as individuals differ in their properties, responses to the environment, and interactions. The collective motion of such heterogeneous systems has been extensively studied, revealing complexities other than those observed in homogeneous systems. Some significant studies have delved into macroscopic colony growth and spatial mixing between populations Deforet et al. (2019); Levis et al. (2019); Liebchen and Levis (2017); Levis and Liebchen (2019); Ventejou et al. (2021); Mecke et al. (2024).
@@
In this study, we investigate an inhomogeneous system focusing on the collective dynamics of a mixed binary mixture of self-propelled particles with different chiralities. We introduce two types of chiral particles, each constituting half of the total population, with one group exhibiting clockwise $(CW)$ rotational tendency and the other anticlockwise $(ACW)$ tendency. This inherent rotation creates heterogeneity within the system. Our primary aim is to understand how these particles interact when they encounter those with particles of opposite nature, particularly under varying parameters like: self-propulsion speed ($v_{0}$) and chirality ($\Omega$). At low $\Omega$, the system shows mixed flocking behavior where particles exhibit only slight rotational tendencies. As $\Omega$ increases, $CW$ and $ACW$ particles begin to separate, forming dense clusters aligned with their rotational directions, resulting in two distinct clusters at moderate $\Omega$ values. Interestingly, at large chirality, particles remain homogeneously mixed at low self-propulsion speeds ($v_{0}=0.001$), rotating uniformly on their axes rather than separating. This phenomenon occurs because as $\Omega$ increases, particles show biased behavior based on their orientation, positive and negative chiral particles tend to move in $CW$ and $ACW$ trajectories, but the self-propulsion speed is not strong enough to separate them. The kinetics of phase separation shows the algebraic growth with growth exponent $1/4$, the same as for other active systems Pattanayak et al. (2021, 2020). This study highlights the intricate interplay between self-propulsion speed and rotational bias due to chirality, revealing various complex behaviors in chiral active systems.
@@
The rest of the article is structured as follows: Section II introduces the model for the dual chiral active particles system. The results and discussion of our system are detailed in Section III. Section III.2 explores the dynamic phases of particles. Finally, Section IV contains the discussion of our results and conclusions.
@@ II Model and Numerical Details
Our system comprises two types of chiral particles, each containing an equal number of $N_{1}=N_{2}=N/2$ particles. These particles possess the same magnitude of chirality with opposite signs. They are distributed across a two-dimensional, off-lattice plane within an $L\times L$ square box with periodic boundaries. The particles are represented as point particles and interact within a defined radius denoted as R. Initially, these particles are randomly dispersed in terms of both position $\textbf{r}_{i}$ and direction ($\theta_{i}$ ranging from $[-\pi,\pi]$), each having a constant self-propelled speed $v_{0}$. The temporal update of positional and direction is done at an interval of $\Delta t$. The particle’s position is updated as follows:
@@
In the given equation, ${\textbf{r}_{i}}(t)$ denote the positions of the $i^{th}$ particle at time $t$, where $i$ can run from $i=1,2,.......N$.
Figure 1: (color online) A schematic representation of the model showing two types of chiral particles with their orientation directions: red for $\Omega^{+}$ (clockwise rotation, $CW$) and blue for $\Omega^{-}$ (anticlockwise rotation, $ACW$). The red arrow in the center of the dashed circle is the $i^{th}$ particle and the dotted line represents the radius of the interaction range $R$. The light red arrows show the effect of random noise $\Delta\theta$. The curved arrows with each particle represent the chiral nature of particles.
${\textbf{v}_{i}}(t)\equiv(v_{0}\cos(\theta_{i}),v_{0}\sin(\theta_{i}))$, represents the velocity of $i^{th}$ particle at time $t$. Here, $\theta_{i}(t)$ is its orientation and each particle has a constant velocity magnitude $v_{0}$. The particles interact locally within their interaction radius $R$ as shown in Fig.1. To incorporate the chiral nature of the particles, the direction of the $i^{th}$ particle is updated according to:
@@
In this context, $\langle\sin\theta_{j}(t)\rangle_{iR}$ and $\langle\cos\theta_{j}(t)\rangle_{iR}$ denote the average values of $\sin\theta$ and $\cos\theta$, respectively, for all particles within the interaction radius $R$ of the $i^{th}$ particle. Chirality is introduced using the parameter $\Omega$, where a positive value (denoted as $\Omega^{+}$) signifies $CW$ rotation and a negative value (denoted as $\Omega^{-}$) represents $ACW$ rotation. The magnitude of the chirality is given by $|\Omega^{\pm}|=\Omega$. $\Delta\theta$ represents the angular noise, which is modeled as $\Delta\theta=\eta\xi_{i}(t)$, $\eta$ represents the noise amplitude and $\xi(t)$ is random variable with zero mean $<\xi_{i}(t)>=0$, delta correlated $<\xi_{i}(t_{1})\xi_{j}(t_{2})>=\delta_{ij}\delta(t_{1}-t_{2})$ and uniformly distributed in $[-\pi,\pi]$. This directional update ensures that the particles demonstrate chiral behavior while considering the average direction of their neighbors within the interaction radius. The noise strength $\eta$ is set to $0.2$, so the system is in the ordered state for a zero-chirality $\Omega=0.0$. The time difference for updating particles position and orientation is fixed to $\Delta t=1.0$. The system is studied for $v_{0}\in(0.001,0.5)$ and $\Omega\in(0.001,0.5)$.
@@
Figure 2: (color online) (I) The figure illustrates the snapshots of the system for $L=40$ and at time $t=2^{17}$. The red dots correspond to particles with positive chirality, denoted by $\Omega^{+}$ and the blue dots represent particles with negative chirality $\Omega^{-}$. These particles are distributed across the coordinate plane, visualizing their spatial arrangement within the system. (II) This figure shows the density difference $\Delta\rho$ for the same system parameters as for (I). The color map indicates the range of $\Delta\rho$: blue regions where $\rho^{-}$ exceeds, red regions where $\rho^{+}$ dominates, and white regions are empty regions. Both panels are depicted in the parameter space of chirality $\Omega$, varying from $0.01$ to $0.5$, and self-propulsion velocity $v_{0}$, ranging from $0.001$ to $0.5$.
@@
Additionally, the interaction radius is fixed at $R=1$, and the total particle density is maintained at $\rho=\dfrac{N}{L^{2}}=1.0$ $(N=1600-10000,L=40-100)$, with $CW$ and $ACW$ particles each having density $0.5$. We let the system evolve, starting with the random homogeneous mixed state of both types of particles, and the position and orientation of particles are updated using Eqs. (1), (2). One simulation step is counted once all the $N$ particles, positions and orientations are updated once. The kinetics of the system is studied for total $t=2^{16}$ time steps and steady-state results are obtained by total simulation steps of $2^{17}$, with time average over the last $2^{15}-2^{17}$ times steps. Further, the averaging of data for the kinetics of the system is performed over $200-500$ independent realizations and the steady state results are obtained by doing the ensemble averaging of data over $10$ independent realizations.
@@ III.1 Steady state characteristics
We first show the snapshots of the system in the steady state at time $t=2^{17}$. In Fig. 2(I) we show the location of two types of particles with two different colors (the red and blue colors represent the $CW$ and $ACW$ particles respectively) and Fig. 2(II) shows the local density difference between two types of particles. The local density of both types of particles $\rho^{\pm}({\bf r})$ in the system is obtained by dividing the whole system into coarse-grained square cells of size $0.625\times 0.625$ and counting the density of particles in each cell centered around point ${\bf r}$ in the system. Further, the normalized density difference of two types of particles is defined as $\Delta\rho({\bf r})=\frac{\rho^{+}({\bf r})-\rho^{-}({\bf r})}{\rho^{+}({\bf r})+\rho^{-}({\bf r})}$. In Figs. 2(II) we show the snapshots of the system, for the density difference $\Delta\rho({\bf r})$ shown in the color bar. The white regions in Fig. 2(II) represent the empty regions. The $+ve$ and $-ve$ values of $\Delta\rho$ indicate the majority of particles are of $CW$ and $ACW$ in nature, respectively. In Fig. 3 we show the orientation $\theta_{i}\in(-\pi,\pi)$ of particles at the same time as for Fig. 2. Now we will explain our observations from Fig. 2(I-II) and 3.
Figure 4: The figures (a), (b), (c) and (d) depict the cartoon of the trajectories of one $CW$ (red) and one $ACW$ (blue) chiral active particles. The solid black circle represents the interaction range with a radius $R=1$, within which the chiral particles interact with each other. The dashed lines show the cartoon of the trajectories of two types of particles with arrows on them representing the directions of velocity of particles. The dots on the dashed line is the location of two types of particles at some arbitrary time. (a) $v_{0}=0.5$ and $\Omega=0.01$, the radius of the chiral particle trajectory ($r_{0}=\frac{v_{0}}{\Omega}=50$) is significantly larger than the interaction range. (b) $v_{0}=0.5$ and $\Omega=0.1$, the trajectories are comparable in size to the interaction radius, resulting in the radius of the trajectory ($r_{0}=\frac{v_{0}}{\Omega}=5$). (c) $v_{0}=0.5$ and $\Omega=0.5$, the radius of trajectory $r_{0}=\frac{v_{0}}{\Omega}=1$ is nearly equal to the interaction radius. (d) $v_{0}=0.2$ and $\Omega=0.5$, the particle trajectories have a smaller radius $r_{0}=\frac{v_{0}}{\Omega}=0.4$ than the interaction radius.
@@
In the lower chirality regime $\Omega=0.01$, the particles remain mixed irrespective of their speed as is evident from the density snapshots in Fig.2II(a-d) and the particle snapshots in Fig2(I)(a-d). However, there is a change in global alignment among the particles in each participating cluster, as the self-propulsion velocity is tuned. For smaller magnitudes of self-propulsion, the particles tend to rotate about their own axis before they go beyond their interaction region and this hinders their chance to interact with new neighbors and system remains nearly homogeneous and weak orientational ordering, as can be seen by the presence of multiple colors in the orientation plot for small $v_{0}=0.001$ and $0.01$ in Figs3(a,b). However as $v_{0}$ is increased, self-propulsion dominates over the rotation due to chirality, in turn enhancing interaction between particles, thus giving rise to a global alignment in the system. This is represented by a color in the majority for higher values of $v_{0}=0.1$ and $0.5$ as shown in Fig.3(c-d). But all through this, the lower magnitude in chirality results in very little orientational difference between both the nature of particles and proves to be insufficient to segregate them ultimately giving rise to a mixed phase. The mixed nature of both the particle types can be seen by the mixing up of colors in Fig. 2(I)(c-d) and small density differences $\Delta\rho$ in Fig.2(II)(c-d).
@@
As we subject the system to a slightly higher chirality we observe varying behaviors based on the magnitude of self-propulsion. Now, as the chirality is higher it dominates over $v_{0}$ for lower self-propulsion values. The particles are then devoid of any chances to interact with each other and mostly rotate in small circles and hence remain in a mixed state, also reducing the global ordering in the process. This can be seen in Fig.2(I-II)(e) and Fig.3(e). On increasing $v_{0}$, we see that the particles start forming small rotating clusters ( as seen in Fig.2(I)(f)) and each cluster has its own local ordering. This is because the chirality is strong enough to create small pockets of rotating clusters and the self-propulsion causes the particles to separate from each other. On further increase in self-propulsion, the particles traverse larger steps before eventually deflecting from their trajectory. But the curvature is strong enough to get the particles separated from each other very soon. Subsequently, it is seen that continuing to increase self-propulsion promotes phase separation and ultimately a global phase-separated state appears for a large $v_{0}=0.5$ containing two distinct clusters of two types of particles.

### v2

# From Flocking to Condensation: Collective Dynamics in Binary Chiral Active Matter

@@ Abstract
Many microswimmers are inherently chiral, and this chirality can introduce fascinating behaviors in a collection of microswimmers. The dynamics become even more intriguing when two types of microswimmers with distinct chirality are mixed. Our study examines a mixture of self-propelled particles with opposite chirality, investigating how the system’s characteristics evolve as the magnitude of chirality is varied. In weakly chiral systems, the particles exhibit similar behavior, leading to a globally flocking phase where both types of particles are well-mixed. However, in an intermediate range of chirality, the condensates of different particles are formed as a result of a competition between chirality and self-propulsion. This competition results in interesting phases within the system. We explore the characteristics of these distinct phases in detail, focusing on the roles of self-propulsion speed and chirality.
@@ I Introduction
Active matter is a fascinating area of study Sumpter (2010); Vicsek and Zafeiris (2012); Mishra et al. (2024); Lavi et al. (2024) that focuses on self-propelled entities such as active colloids Bechinger et al. (2016), insects swarming in large groups Cavagna et al. (2017), fish displaying schooling in the ocean Ward et al. (2008), and birds flocking in the sky Ballerini et al. (2008), groups of animals at the macroscopic scale Cavagna and Giardina (2014), etc. These entities move due to internal mechanisms such as chemical reactions, cilia, flagella, and internal motors that utilize the input from the surrounding mediums. Understanding collective behavior in active matter is crucial to interpret collective phenomena. Flocking, which refers to the collective motion of organisms, is also observed in inanimate macroscopic systems, such as active granular rods Kumar et al. (2014). This widespread phenomenon often involves transitions from disordered to long-range ordered states, influenced by factors such as density and noise Vicsek et al. (1995a).
@@
The dynamics of active agents can be categorized into two main types: one involves nearly straight movement with random directional changes, creating a sense of self-propulsion without any perception of rotation, as seen in active colloids like Janus particles, molecular motors Jülicher et al. (1997), natural low-Reynolds-number swimmers Najafi and Golestanian (2004); Ebrahimian et al. (2015); Bennett and Golestanian (2013), active particles on periodic lattice Chamolly et al. (2017) and polar active granular particles Zhang et al. (2017). Particles with a sense of rotation or chiral particles characterize the second type. Systematically, particles can tend to rotate clockwise $(CW)$ or anticlockwise $(ACW)$. The motion is circular in two dimensions and helical in three dimensions Löwen (2016); Friedrich (2016). These chiral particles possess a characteristic self-propulsion velocity and a chirality that determines the rate of directional change in the absence of fluctuations. The radius of their trajectories is determined by the ratio of self-propulsion velocity to chirality Caprini et al. (2024). This type of chiral motion is observed in various systems, including biological circle swimmers such as bacteria DiLuzio et al. (2005), E.coli Berg and Turner (1990), sperm cells Riedel et al. (2005), artificial microswimmers of L shape Kümmel et al. (2013), and magnetotactic bacteria in rotating external fields Ērglis et al. (2007).
@@
Although most of the research on bacterial swarming and collective swimming has focused on homogeneous systems Harshey (2003, 2003); Sokolov et al. (2007); Copeland and Weibel (2009), which involve agents having identical features. The natural world encounters mixtures of particles of different chiralities Schumacher et al. (2017); Jolles et al. (2020); Ao et al. (2015). These differences arise from variations in their size Peled et al. (2021), diverse orientations, and behavior of mixed species Benisty et al. (2015); Zuo and Wu (2020). Complex systems are naturally diverse, as individuals vary in their characteristics, environmental reactions, and interactions. The collective behavior of such diverse systems has been thoroughly examined, revealing complexities that differ from those found in uniform systems.
Recent experiments have explored mixtures of two types of chiral particles, revealing fascinating collective behaviors. In some studies, chiral particles with clockwise $(CW)$ and anticlockwise $(ACW)$ rotations are generated using confined geometries with polarized walls Barois et al. (2020). In the case of run-and-tumble particles, a high tumbling rate can induce chiral motion, causing bacteria to behave like chiral swimmers Ariel et al. (2018). This phenomenon also demonstrates a method for sorting particles based on their physical properties using asymmetric barriers without external driving forces. Reichhardt and Reichhardt (2013). In one of our previous works Eswaran and Mishra (2024), we also found the synchronized rotation of active Brownian particles due to chemotaxis, which is another way of introducing the chirality in the dynamics of particles due to the medium. Reversible pattern formation and edge transport have also been studied in mixtures of active chiral and passive disks, where the degree of segregation depends on orbit radius and density Reichhardt and Reichhardt (2019a). Additionally, probe particles driven through chiral fluids exhibit transverse Hall-like motion, revealing the role of chirality in microrheological response and jamming behavior Reichhardt and Reichhardt (2019b). Some significant studies have delved into macroscopic colony growth and spatial mixing between populations Deforet et al. (2019); Levis et al. (2019); Liebchen and Levis (2017); Levis and Liebchen (2019); Ventejou et al. (2021); Mecke et al. (2024).
@@
In this study, we consider a minimal model of point-like chiral active particles interacting through local velocity alignment in a binary mixture with equal and opposite chirality. Our model is inspired by the Vicsek modelVicsek et al. (1995b), which introduced alignment-based interactions among self-propelled particles as a paradigm for collective motion. However, unlike the Vicsek model—which considers homogeneous particles aligning their velocity—we introduce heterogeneity through a binary mixture of oppositely chiral particles. We introduce heterogeneity through a binary mixture of oppositely chiral particles. We introduce two types of chiral particles, each constituting half of the total population, with one group exhibiting clockwise $(CW)$ rotational tendency and the other anticlockwise $(ACW)$ tendency. This inherent rotation creates heterogeneity within the system. Our primary aim is to understand how these particles interact when they encounter those with particles of opposite nature, particularly under varying parameters like: self-propulsion speed ($v_{0}$) and chirality ($\Omega$). At low $\Omega$, the system shows mixed flocking behavior where particles exhibit only slight rotational tendencies. As $\Omega$ increases, $CW$ and $ACW$ particles separate, forming dense clusters aligned with their rotational directions, resulting in two distinct clusters at moderate $\Omega$ values. Interestingly, particles remain homogeneously mixed at large chirality at low self-propulsion speeds ($v_{0}=0.001$), rotating uniformly on their axes rather than separating. This phenomenon occurs because, as $\Omega$ increases, particles show biased behavior based on their orientation; positive and negative chiral particles tend to move in $CW$ and $ACW$ trajectories, but the self-propulsion speed is not strong enough to separate them. The self-sorting of the chiral species in the mixture of chiral species is also observed in the experiment of Gübitz and Schmid (2001), where the active particles segregate based on chirality without the need for chemical reactions. The clusters of segregated chiral particles of the same type observed in our study resemble the structures seen in experiments Pi et al. (2022); Sirota-Madi et al. (2010); Denk et al. (2016). The phase separation kinetics show the algebraic growth with a growth exponent close to $1/4$. The same as for other active systems Pattanayak et al. (2021, 2020); Wittkowski et al. (2014). This study highlights the intricate interplay between self-propulsion speed and rotational bias due to chirality, revealing various complex behaviors in chiral active systems.
@@
The rest of the article is structured as follows: Section II introduces the model for the dual chiral active particle system. The results and discussion of our system are detailed in Section III. Section III.2 explores the dynamic phases of particles. Finally, Section IV discusses our results and conclusions.
@@ II Model and Numerical Details
Our system comprises two chiral particles, each containing an equal number of $N_{1}=N_{2}=N/2$ particles. These particles possess the same magnitude of chirality with opposite signs. They are distributed across a two-dimensional, off-lattice plane within an $L\times L$ square box with periodic boundaries. The particles are represented as point particles and interact within a defined radius denoted as R. Initially, these particles are randomly dispersed in terms of both position $\textbf{r}_{i}$ and direction ($\theta_{i}$ ranging from $\pi,\pi]$), each having a constant self-propelled speed $v_{0}$. The temporal update of position and direction is done at an interval of $\Delta t$. The particle’s position is updated as follows:
@@
In the given equation, ${\textbf{r}_{i}}(t)$ denotes the positions of the $i^{th}$ particle at time $t$, where $i$ can run from i=1,2,.Ni=1,2,.......N.
Figure 1: (color online) A schematic representation of the model showing two types of chiral particles with their orientation directions: red for $\Omega^{+}$ (clockwise rotation, CWCW) and blue for $\Omega^{-}$ (anticlockwise rotation, ACWACW). The red arrow in the center of the dashed circle is the $i^{th}$ particle, and the dotted line represents the radius of the interaction range $R$. The light red arrows show the effect of random noise $\Delta\theta$. The curved arrows with each particle represent the chiral nature of particles.
${\textbf{v}_{i}}(t)\equiv(v_{0}\cos(\theta_{i}),v_{0}\sin(\theta_{i}))$, represents the velocity of an $i^{th}$ particle at time $t$. Here, $\theta_{i}(t)$ is its orientation, and each particle moves with a constant self-propulsion speed $v_{0}$. The particles interact locally within their interaction radius $R$ as shown in Fig.1. To incorporate the chiral nature of the particles, the direction of the $i^{th}$ particle is updated according to:
@@
In this context, $\langle\sin\theta_{j}(t)\rangle_{iR}$ and $\langle\cos\theta_{j}(t)\rangle_{iR}$ denote the average values of $\sin\theta$ and $\cos\theta$, respectively, for all particles within the interaction radius $R$ of the $i^{th}$ particle. Chirality is introduced using the parameter $\Omega$, where a positive value (denoted as $\Omega^{+}$) signifies $CW$ rotation and a negative value (denoted as $\Omega^{-}$) represents $ACW$ rotation. The magnitude of the chirality is given by $|\Omega^{\pm}|=\Omega$. $\Delta\theta$ represents the angular noise, which is modeled as $\Delta\theta=\eta\xi_{i}(t)$, $\eta$ represents the noise amplitude and $\xi(t)$ is a random variable with zero mean $\xi_{i}(t)>=0$, delta correlated $\xi_{i}(t_{1})\xi_{j}(t_{2})>=\delta_{ij}\delta(t_{1}-t_{2})$ and uniformly distributed in $\pi,\pi]$. This directional update ensures that the particles demonstrate chiral behavior while considering the average direction of their neighbors within the interaction radius. The noise strength $\eta$ is set to $0.2$, so the system is in the ordered state for a zero chirality $\Omega=0.0$. The time difference for updating particles’ position and orientation is fixed to $\Delta t=1.0$. The system is studied for $v_{0}\in(0.001,0.8)$ and $\Omega\in(0.001,0.5)$.
@@
Figure 2: (color online) (I) The figure illustrates the snapshots of the system for $L=40$ and at time $t=2^{17}$. The red dots correspond to particles with positive chirality, denoted by $\Omega^{+}$, and the blue dots represent particles with negative chirality $\Omega^{-}$. These particles are distributed across the coordinate plane, visualizing their spatial arrangement. (II) This figure shows the density difference $\Delta\rho$ for the same system parameters as for (I). The color map indicates the $\Delta\rho$ range: blue regions where $\rho^{-}$ exceeds, red regions where $\rho^{+}$ dominates, and white regions are empty. Both panels are depicted in the parameter space of chirality $\Omega$, varying from $0.01$ to $0.5$, and self-propulsion velocity $v_{0}$, ranging from $0.001$ to $0.5$.
@@
Additionally, the interaction radius is fixed at $R=1$, and the total particle density is maintained at $\rho=\dfrac{N}{L^{2}}=1.0$ (N=40010000,L=20100)(N=400-10000,L=20-100), with $CW$ and $ACW$ particles each having density $0.5$. We let the system evolve, starting with the random homogeneous mixed state of both types of particles, and the position and orientation of particles are updated using equations (1), (2). Once all the $N$ particles’ positions and orientations are updated, one simulation step is counted. The kinetics of the system are studied for a total of $t=2^{16}$ time steps, and steady-state results are obtained by total simulation steps of $2^{17}$, with time average over the last $2^{15}-2^{17}$ time steps. Further, data averaging for the system’s kinetics over 200500200-500 independent realizations is performed. The steady-state results are obtained by ensemble averaging data over $10$ independent realizations.
@@ III.1 Steady state characteristics
We first show the snapshots of the system in the steady state at time $t=2^{17}$. In Fig.2(I), we show the location of two types of particles with two different colors (the red and blue colors represent the $CW$ and $ACW$ particles, respectively), and Fig.2(II) shows the local density difference between the two types of particles. The local density of both types of particles $\rho^{\pm}({\bf r})$ in the system is obtained by dividing the whole system into coarse-grained square cells of size $0.625\times 0.625$ and counting the density of particles in each cell centered around point $\bf r$} in the system. Further, the normalized density difference of the two types of particles is defined as $\Delta\rho({\bf r})=\frac{\rho^{+}({\bf r})-\rho^{-}({\bf r})}{\rho^{+}({\bf r})+\rho^{-}({\bf r})}$. In Fig.2(II), we show the snapshots of the system for the density difference $\Delta\rho({\bf r})$ shown in the color bar. The white regions in Fig.2(II) represent the empty regions. The positive and negative values of $\Delta\rho$ indicate that most particles are of $CW$ and $ACW$ in nature, respectively. In Fig.3 we show the orientation $\theta_{i}\in(-\pi,\pi)$ of particles at the same time as for Fig.2. Now we will explain our observations from Figs.2(I-II) and 3.
@@
In the lower chirality regime $\Omega=0.01$, the particles remain mixed irrespective of their speed, as is evident from the density snapshots in Figs.2II(a-d) and the particle snapshots in Figs.2(I)(a-d). However, there is a change in global alignment among the particles in each participating cluster as the self-propulsion velocity is tuned. For smaller magnitudes of self-propulsion, the particles rotate about their axis before they go beyond their interaction region, which hinders their chance to interact with new neighbors. The system remains nearly homogeneous and has weak orientational ordering, as can be seen by multiple colors in the orientation plot for small $v_{0}=0.001$ and $0.01$ in Figs.3(a,b). However, as $v_{0}$ increases, self-propulsion dominates the rotation due to chirality, enhancing particle interaction and giving rise to a global alignment in the system. This is represented by a color in the majority for higher values of $v_{0}=0.1$ and $0.5$ as shown in Figs.3(c-d). However, all through this, the lower magnitude in chirality results in very little orientational difference between the nature of particles and proves insufficient to segregate them, ultimately giving rise to a mixed phase. The mixing can be seen in the mixed nature of both the particle types of colors in Figs.2(I)(c-d) and minor density differences $\Delta\rho$ in Figs.2(II)(c-d).
@@
As we subject the system to a slightly higher chirality, we observe varying behaviors based on the magnitude of self-propulsion. As the chirality is higher, it dominates over $v_{0}$ for lower self-propulsion values. The particles are then devoid of any chances to interact with each other, mostly rotating in small circles, and hence remain mixed, reducing the global ordering in the process. This can be seen in Figs.2(I-II)(e) and Fig.3(e). On increasing $v_{0}$, we see that the particles start forming small rotating clusters (as seen in Fig.2(I)(f)), and each cluster has its local ordering. This is because the chirality is strong enough to create small pockets of rotating clusters, and the self-propulsion causes the particles to separate. On further increase in self-propulsion, the particles traverse larger steps before eventually deflecting from their trajectory, the curvature is strong enough to separate the particles in a short period.
Figure 4: The figures (a), (b), (c), and (d) depict the cartoon of the trajectories of one CWCW (red) and one ACWACW (blue) chiral active particles. The solid black circle represents the interaction range with a radius $R=1$, within which the chiral particles interact. The dashed lines show the cartoon of the trajectories of two types of particles, with arrows representing the directions of the velocity of the particles. The dots on the dashed line are the locations of two types of particles at some arbitrary time. (a) $v_{0}=0.5$ and $\Omega=0.01$, the radius of the chiral particle trajectory ($r_{0}=\frac{v_{0}}{\Omega}=50$) is significantly larger than the interaction range. (b) $v_{0}=0.5$ and $\Omega=0.1$, the trajectories are comparable in size to the interaction radius, resulting in the radius of the trajectory ($r_{0}=\frac{v_{0}}{\Omega}=5$). (c) $v_{0}=0.5$ and $\Omega=0.5$, the radius of trajectory $r_{0}=\frac{v_{0}}{\Omega}=1$ is nearly equal to the interaction radius. (d) $v_{0}=0.2$ and $\Omega=0.5$, the particle trajectories have a smaller radius $r_{0}=\frac{v_{0}}{\Omega}=0.4$ than the interaction radius.
From the above discussion, we can conclude that at intermediate values of $(\Omega)$, when the two types of particles undergo phase segregation, the resulting clusters behave more like condensation than traditional phase separation. Notably, as the system size $(L)$ increases, the size of the clusters $l_{sat}$ remains roughly constant, Figs.5(b,c), while their density rises according to a power-law scaling Fig.5(a). Specifically, the density difference in condensed clusters is as $(\Delta\rho^{*}\sim L^{2})$ where $\Delta\rho^{*}=\{\rho^{+}({\bf r})-\rho^{-}({\bf r})\}_{max}$, where $\{...\}_{max}$ means the maximum of local density difference $\rho^{+}$ and $\rho^{-}$. Further, the $\Delta\rho^{*}$ is obtained by averaging over time in the steady state and over $500$ independent realizations. The variation of $\Delta\rho^{*}$ as $L^{2}$ means that the number of particles in each cluster grows much faster than linear dependence, leading to denser structures without an increase in cluster size. In contrast, conventional phase segregation typically involves coarsening, where the cluster size increases with the system size. In our case, the saturation length of the clusters does not change significantly, indicating that the clusters maintain a finite size regardless of the growing system size. Instead, additional particles condense into existing clusters, rapidly increasing density. This behavior aligns with condensate clusters, where a portion of the system’s mass condenses into high-density regions, as seen in other active matter systems Wang et al. (2024). Thus, these phase-separated structures are better classified as condensation-driven clusters.